# Symbolic Mixture-of-Experts: Adaptive Skill-based Routing for Heterogeneous Reasoning

### v2

@@ 4.3 Analysis
Symbolic-MoE profiles models based on their skills and leverages this information to recruit experts effectively. To underscore the importance of this step, we compare several alternative selection strategies in Table 4, evaluating accuracy on GPQA. Specifically, we assess performance when selecting the top-$k$ agents overall and when choosing $k$ agents at random. In the top-$k$ approach, experts are fixed as the best-performing models for the task, whereas in the random-$k$ strategy, the selected experts vary across instances. Our results demonstrate that skill-based selection is essential. Notably, although selecting the top $k$ experts for a task may seem intuitive, it consistently underperforms compared to Symbolic-MoE’s adaptive instance-level expert selection. Interestingly, top-$5$ selection performs worse than top-$3$ selection, suggesting that a broader selection increases the likelihood of including weaker models, leading to performance degradation. Additionally, the random selection strategy consistently harms performance, showing a $12.86\%-15.61\%$ drop compared to Symbolic-MoE, likely also due to the inclusion of weaker experts. Overall, skill-based instance-level selection consistently outperforms more coarse-grained approaches, with Symbolic-MoE surpassing MoA by $4.92\%$ as also shown in Table 1. These findings emphasize the value of leveraging performance-based statistics and individual skill assessments for expert selection.
@@
Role and Selection of the Aggregator.
Unlike most of our discussion-based multi-agent baselines, Symbolic-MoE collects a single CoT and answer from each expert and combines them via an aggregator. Besides the efficiency gain as shown in Table 2, we investigate the role the aggregator plays in our framework. While experts are selected per instance, the aggregator is chosen based at the task, as we find that reasoning ability does not necessarily translate to effective aggregation. Table 5 examines the impact of aggregator selection in Symbolic-MoE, comparing three strategies: (1) a randomly chosen aggregator from the model pool, (2) an instance-specific top-1 aggregator based on model profiling, and (3) a task-specific aggregator determined by task-level performance. Evaluated on MMLU-Pro and GPQA, the results indicate that a random aggregator substantially degrades performance, indicating that the aggregator plays a crucial role. While the instance-specific top-1 aggregator improves outcomes on both datasets, the task-specific aggregator outperforms it on MMLU-Pro and performs comparably on GPQA. We further find that the similar performance of instance-specific and task-specific aggregation on GPQA is due to a high degree of overlap in selected aggregators. Overall, these findings suggest that a model being a good reasoner does not imply it will be a good aggregator, supporting our decision to use task-based aggregation.
@@ 5 Discussion and Conclusion
A key feature highlighted in Table 1 is the consistency of Symbolic-MoE’s performance. While baseline methods occasionally do well in isolated settings (e.g. MoA on MMLU-Pro, ReConcile on MedMCQA) it is important to highlight that no baseline does well consistently across settings. This means that -- without Symbolic-MoE -- getting a strong overall result would require evaluating all the baseline methods and choosing the best settings manually. In contrast to the baselines, Symbolic-MoE consistently achieves high performance without human intervention. By automatically selecting discussion agents, Symbolic-MoE provides a consistent recipe that generalizes across domains.
Modularity.
@@
Another key advantage of Symbolic-MoE is its modularity. Unlike typical Mixture-of-Experts (MoEs) frameworks, which need to be trained end-to-end from scratch in a centralized manner, Symbolic-MoE uses the symbolic output channel of existing models to combine experts. This gradient-free approach enables seamless integration of pre-trained models without updates, allowing them to be trained independently and distributedly. Such delegation enhances domain specialization for each model. Moreover, while standard MoEs have a fixed size determined before training, Symbolic-MoE can adaptively grow and evolve as models are updated. Given the rapid advancements in LLMs, cost-effective and efficient updates are essential -- state-of-the-art models are often replaced within months. Symbolic-MoE’s modular and gradient-free design simplifies these updates, requiring only a few calls to obtain a new model’s agent profile.
Connections to Inference-Time Scaling.
@@
Like other LLM-discussion frameworks, Symbolic-MoE can be seen as a form of multi-model inference-time scaling. Past work (Wang et al., 2023a; Snell et al., 2024; Shao et al., 2024) has highlighted the benefits of adding inference-time compute, ranging from simply sampling multiple responses (as done in our SC baselines) to more complex strategies such as refinement (Madaan et al., 2023a; Chen et al., 2024b). Symbolic-MoE surpasses several such baselines (Table 1) by adaptively selecting the optimal set of expert models while avoiding the costly discussion process, reducing overhead while improving performance. Moreover, our novel batching method enables Symbolic-MoE to run efficiently on a single GPU while remaining flexible for acceleration with multiple GPUs.
Conclusion.
@@
We introduced Symbolic-MoE, a scalable MoE framework that combines models through their symbolic output (i.e., via natural language discussion). Symbolic-MoE infers which skills are needed for a given problem and recruits agents based on those skills to engage in a discussion about a given input, guiding the discussion to reach a better consensus. On four diverse reasoning datasets, Symbolic-MoE outperforms standard inference-time scaling methods as well as other debate frameworks and other mixture-of-agents methods, leading to consistently strong performance across domains without human intervention. Symbolic-MoE’s average performance across heterogeneous tasks is in fact stronger than that of advanced proprietary models such as GPT4o-mini. Moreover, unlike past work that requires multiple models to be loaded on separate GPUs running in parallel, Symbolic-MoE introduces a novel batching strategy that allows us to run on a single GPU in roughly the same amount of time, obtaining the best of both worlds in terms of performance and efficiency.
@@
Symbolic-MoE also relies on skills inferred from a small validation set to set the agent profiles. In our experiments, we ensure fair comparisons to the baselines by choosing models for the baselines according to the same validation set, giving the baselines equal access to the data. Skill inference relies on the Keyword LLM being sufficiently trained on a given domain to infer relevant skills -- empirically, we find that this is the case across a variety of domains. Overall, Symbolic-MoE will continue to improve with better skill inference modules, which can easily be swapped in.
@@ Acknowledgements
Appendix A Dataset Statistics and Licenses
Validation Size Test Size License
MMLU-Pro (Wang et al., 2024c) 350 2,100 Apache License
AIME (MAA, 2024) 354 30 CC0
GPQA (Rein et al., 2023) 249 198 MIT License
MedMCQA (Pal et al., 2022) 504 4,183 MIT License
Appendix B Model Pool
@@
Appendix C Performance on the Validation Set
@@
Appendix D Performance of Each Model as an Aggregator
@@
Appendix E Distribution of Experts
We present the distribution of recruited experts across different datasets in Fig. 4. As noted in Section 3.2.1, we trim experts with occurrences below 5% to reduce model loading time. This is visualized in Fig. 4, where the top row shows the distribution before trimming, and the bottom row shows the distribution after trimming. The distribution varies significantly across datasets -- on more diverse datasets such as MMLU-Pro, the recruited experts are also more varied. In contrast, for AIME and GPQA, which focus more on math and science, the recruited experts are dominated by a few models.
Appendix F Algorithm
@@ Appendix G Prompts
Follow this format: Keywords: <keyword1>, <keyword2>...
@@
Provide your step-by-step reasoning first, and then print ‘‘The answer is (X)’’ where X is the answer choice (one capital letter), at the end of your response.
@@
Provide your step-by-step reasoning first, and then print ‘‘The answer is (X)’’ where X is the answer choice (one capital letter), at the end of your response.
Appendix H Symbolic-MoE as a Sparse Mixture-of-Expert

### v3

@@ 4.4 Additional Analysis
Utility of the Model Profile. Symbolic-MoE profiles models based on their skills and leverages this information to recruit experts effectively. To underscore the importance of this step, we compare several alternative selection strategies in Table 4, evaluating accuracy on GPQA. Specifically, we assess performance when selecting the top-$k$ agents overall and when choosing $k$ agents at random. In the top-$k$ approach, experts are fixed as the best-performing models for the task, whereas in the random-$k$ strategy, the selected experts vary across instances. Our results demonstrate that skill-based selection is essential. Notably, although selecting the top $k$ experts for a task may seem intuitive, it consistently underperforms compared to Symbolic-MoE’s adaptive instance-level expert selection. Interestingly, top-$5$ selection performs worse than top-$3$ selection, suggesting that a broader selection increases the likelihood of including weaker models, leading to performance degradation. Additionally, the random selection strategy consistently harms performance, showing a $12.86\%$ to $15.61\%$ drop compared to Symbolic-MoE, likely also due to the inclusion of weaker experts. Overall, skill-based instance-level selection consistently outperforms more coarse-grained approaches, with Symbolic-MoE surpassing MoA by $4.92\%$ as also shown in Table 1.
Model MATH-Hard Omni
Debate 83.31 42.93
MoA 86.48 47.36
Self-MoA 88.75 48.75
ReConcile 79.31 42.62
Symbolic-MoE 91.62 52.03
Table 5: Evaluation on MATH-Hard and OmniMATH using profiles from AIME.
Symbolic-MoE generalizes to unseen tasks. Given the constant introduction of new, unseen data in LLM deployments, we further assess whether Symbolic-MoE can generalize to unseen settings using existing model profiles on two additional benchmarks: OmniMATH (Gao et al., 2025) and MATH Hard (Hendrycks et al., 2021b). OmniMATH is an Olympiad-level math benchmark consisting of 4.4k test samples, while MATH Hard is the level 5 (the most challenging) subset of MATH, comprising 1.3k problems. For these benchmarks, we did not create model profiles using their respective validation sets. Instead, we reused the model profiles obtained from AIME. As for the multi-agent baselines, we also use the top-3 models from AIME. Results in Table 5 show that Symbolic-MoE generalizes effectively, even when evaluated on unseen tasks and with model profiles not tailored to the test domain. Notably, the transferred Symbolic-MoE outperforms the best baseline (Self-MoA) by 2.87% on MATH-Hard and 3.28% on OmniMATH.
@@
Role and selection of the aggregator. Unlike most of our discussion-based multi-agent baselines, Symbolic-MoE collects a single CoT and answer from each expert and combines them via an aggregator. This provides efficiency gains, as shown in Table 2; here, we investigate the role of aggregator in our framework. While experts are selected per instance, the aggregator is chosen per task, as we find that reasoning ability does not necessarily translate to effective aggregation. Table 6 examines the impact of aggregator selection in Symbolic-MoE, comparing three strategies: (1) a randomly chosen aggregator from the model pool (Random), (2) an instance-level aggregator selection based on model profiling (Adaptive), and (3) a task-specific aggregator determined by task-level performance (Task-specific). Evaluated on MMLU-Pro and GPQA, the results indicate that a random aggregator substantially degrades performance, showing that the aggregator plays a crucial role. While the instance-level aggregator improves outcomes on both datasets, the task-specific aggregator outperforms it on MMLU-Pro and performs comparably on GPQA. We further find that the similar performance of instance-specific and task-specific aggregation on GPQA is due to a high degree of overlap in selected aggregators. Overall, this suggests that good reasoners will not always be good aggregators, supporting our task-based selection.
Expert Aggregator GPQA
Random Task-Specific 31.82
Recruited Random 51.52
Recruited Majority Vote 53.54
Recruited Task-Specific 57.78
Table 7: Accuracy on GPQA using different expert selection and aggregation strategies.
Synergy between expert and aggregator selection. As demonstated that the selection of an aggregator plays an important role, we further investigate the synergy between good aggregator selection and good expert selection. In Table 7, we experiment with two expert selection strategies: (1) randomly recruiting $k$ experts without using model profiles, and (2) using our model profiles to recruit experts. We also employ three settings for aggregator selection: (1) based on the task performance as Symbolic-MoE uses, (2) randomly selecting an aggregator, and (3) use majority vote to select the final answer without using any aggregator. Our findings indicate that the combination of strong models with a task-specific aggregator yields the highest performance. When the aggregator is suboptimal, majority voting can serve as a robust alternative. However, when the expert models themselves are weak (chosen randomly), even a strong aggregator cannot compensate for the performance drop.
Keyword LLM MMLU-Pro GPQA
Llama 3.1 8B 64.19 56.62
Gemma 2 9B 64.02 57.01
Qwen 2.5 7B 63.71 57.78
Table 8: Keyword LLM has little effect on the final performance.
Sensitivity to the keyword LLM. We choose Qwen 2.5 7B (Qwen Team, 2024) as the “Keyword LLM” to generate the required skills for each instance during both preprocessing and inference. Here, we investigate the sensitivity of the results to the choice of the Keyword LLM, testing three different models: Qwen 2.5 7B (Qwen Team, 2024), Llama 3.1 8B (AI, 2024), and Gemma 2 9B (Team et al., 2024b). As shown in Table 8, the final performance remains consistent regardless of the chosen model, indicating that the selection of the Keyword LLM has minimal influence on performance.
@@ 5 Discussion and Conclusion
A key feature highlighted in Table 1 is the consistency of Symbolic-MoE’s performance. While baseline methods occasionally do well in isolated settings (e.g. MoA on MMLU-Pro, ReConcile on MedMCQA) it is important to highlight that no baseline does well consistently across settings. This means that – without Symbolic-MoE – getting a strong overall result would require evaluating all the baseline methods and choosing the best settings manually. In contrast, Symbolic-MoE achieves high performance without human intervention. By automatically recruiting the experts based on skills needed for each instance, Symbolic-MoE provides a robust recipe that generalizes across domains.
@@
Modularity. Another key advantage of Symbolic-MoE is its modularity. Unlike typical Mixture-of-Experts (MoE) frameworks, which need to be trained end-to-end from scratch in a centralized manner, Symbolic-MoE uses the symbolic output channel of existing models to combine experts. This gradient-free approach enables seamless integration of pre-trained models without updates, allowing them to be trained independently and distributedly. Such delegation enhances domain specialization for each model. Moreover, while standard MoEs have a fixed size determined before training, Symbolic-MoE can adaptively grow and evolve as models are updated. Given the rapid advancements in LLMs, cost-effective and efficient updates are essential – state-of-the-art models are often replaced within months. Symbolic-MoE’s modular and gradient-free design simplifies the incorporation of these updates, requiring only a few calls to obtain a new model’s profile, which can then be easily plugged into this framework. It is also straightforward to increase the number of experts recruited at test time, which is usually fixed in the typical MoE setting.
@@
Connections to Inference-Time Scaling. Like other LLM-discussion frameworks, Symbolic-MoE can be seen as a form of multi-model inference-time scaling. Past works (Snell et al., 2024; Shao et al., 2024) have highlighted the benefits of adding inference-time compute, ranging from simply sampling multiple responses (as in our SC baselines) to more complex strategies such as refinement (Madaan et al., 2023a; Chen et al., 2024b). Symbolic-MoE surpasses several such baselines (Table 1) by adaptively selecting the optimal set of expert models while avoiding the costly discussion process, reducing overhead while improving performance. Moreover, our novel batching method enables Symbolic-MoE to run efficiently on a single GPU and flexibly scale across multiple GPUs.
@@
Conclusion. We introduced Symbolic-MoE, a scalable MoE framework that combines models through their symbolic outputs (i.e., via natural language discussion). Symbolic-MoE infers which skills are needed for a given problem and recruits agents accordingly to engage in a discussion about a given input, guiding the discussion to reach a better consensus. On four diverse reasoning datasets, Symbolic-MoE outperforms standard inference-time scaling methods, debate frameworks and recent mixture-of-agents methods, leading to consistently strong performance across domains without human intervention. Symbolic-MoE’s average performance across heterogeneous tasks is in fact stronger than that of advanced proprietary models such as GPT4o-mini. Moreover, unlike past work that requires multiple models to be loaded on separate GPUs running in parallel, Symbolic-MoE introduces a novel batching strategy that allows us to run on a single GPU in roughly the same amount of time, obtaining the best of both worlds in terms of performance and efficiency.
@@
Symbolic-MoE also relies on skills inferred from a small validation set to set the agent profiles. In our experiments, we ensure fair comparisons to the baselines by choosing models for the baselines according to the same validation set, giving the baselines equal access to the data. Skill inference relies on the Keyword LLM being sufficiently trained on a given domain to infer relevant skills – empirically, we find that this is the case across a variety of domains. Overall, Symbolic-MoE will continue to improve with better skill inference modules, which can easily be swapped in.
@@ Acknowledgements
Appendix A Illustrations of Variations in Batch Inference
As discussed in Section 3.3, there are mainly three different ways to serve multiple LLMs to solve every instance adaptively. We illustrate these variations in Fig. 3. Our batch inference method allows for the speedups featured in Table 2.
Appendix B Baselines Details
We compare against four categories of baselines.
  • Zero-shot single-model methods: This category includes proprietary models such as GPT-4o-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024a), and DeepSeek-V3 (DeepSeek-AI et al., 2025b); high-capacity open-source models like Qwen2.5 72B (Qwen et al., 2025) and Llama 3.3 70B (AI@Meta, 2024); and strong distilled 7B models such as QwenR1 (DeepSeek-AI et al., 2025a). For reference, we also report the best task-specific model from our pool for each task, denoted as Task-Best.
Appendix C Implementation Details
We conduct our experiments for Symbolic-MoE and other single-model baselines on a single A6000 GPU with 48 GB of memory, while MoA and ReConcile are executed on 8 A6000 GPUs for parallelization. For the 70B models, we use the original version without quantization and perform inference on 4 A6000 GPUs. All open-source models utilize vLLM (Kwon et al., 2023) for inference. The temperature is set to 0.7 for all methods. The maximum output token length is fixed at $4096$ for all models, except for QwenR1 and LlamaR1, which have a limit of $32768$ since they are trained with longer trajectories and tend to generate longer outputs. All results, except those from proprietary models (due to budget constraints), are averaged over three random seeds. Further details on the model pool, distribution of the expert recruited, and all the prompts we use can be found in Table 9 and Appendix L.
Appendix D Model Pool
We provide the full list of our model pool in Table 9, including their names, sizes, and publicly available checkpoints on Huggingface. Most of the model sizes are 7 to 8 billion.
@@ Appendix D Model Pool
Appendix E Performance on the Validation Set
@@
Appendix F Performance of Each Model as an Aggregator
@@
Appendix G Distribution of Experts
We present the distribution of recruited experts across different datasets in Fig. 4. As noted in Section 3.2.1, we trim experts with occurrences below 5% to reduce model loading time. In Fig. 4, the top row shows the distribution before trimming, and the bottom row shows the distribution after trimming. The distribution varies significantly across datasets – on more diverse datasets such as MMLU-Pro, the recruited experts are also more varied. In contrast, for AIME and GPQA, which focus more on math and science, the recruited experts are dominated by a few models.
@@ Appendix G Distribution of Experts
Appendix H Test-time Token Count Analysis
In addition to measuring GPU run time in Table 2, we compare the test-time token count with multi-agent baselines. As shown in Fig. 5, Symbolic-MoE uses fewer tokens than Self-MoA while achieving a significant performance gain. However, compared to MoA and ReConcile, Symbolic-MoE generates more tokens, particularly on GPQA. The primary reason, as illustrated in Fig. 4, is that Symbolic-MoE predominantly recruits LlamaR1 and QwenR1, both of which are trained with long reasoning trajectories, resulting in substantially longer outputs compared to other models. This explains why Symbolic-MoE requires less GPU run time despite producing more tokens: by skipping the expensive multi-round discussions, we eliminate the time spent loading and offloading models. However, the inherent verbosity of the R1 models contributes to the higher token count.
Figure 5: Comparison of the test-time token count used in different methods.
MMLU-Pro GPQA
Acc $\uparrow$ Time $\downarrow$ Acc $\uparrow$ Time $\downarrow$
w/o Trimming 63.94 18.83 55.26 21.78
w/ Trimming 63.71 12.27 57.78 10.85
Table 12: Trimming the low-frequent experts improves both accuracy and efficiency.
Appendix I The Effect of Trimming and Re-sampling
As described in Section 3.3.1, we trim the recruited experts if their occurrence falls below $5\%$ of the total selections. Here, we analyze the impact of this trimming process. Without trimming, the diversity of selected experts increases, but the model loading time also increases. Table 12 presents a quantitative comparison of accuracy and GPU run time using 4 GPUs. As expected, trimming reduces GPU run time across both datasets by minimizing the need to load infrequently used models. Interestingly, we also observe that trimming improves accuracy on GPQA. This improvement may be due to the fact that after trimming, only LlamaR1 and QwenR1 remain as the recruited experts, which are particularly effective on this task.
Appendix J Dataset Statistics and Licenses
Validation Size Test Size License
MMLU-Pro (Wang et al., 2024c) 350 2,100 Apache License
AIME (MAA, 2024) 354 30 CC0
GPQA (Rein et al., 2023) 249 198 MIT License
MedMCQA (Pal et al., 2022) 504 4,183 MIT License
Appendix K Algorithm
@@ Appendix I The Effect of Trimming and Re-sampling
Follow this format: Keywords: <keyword1>, <keyword2>…
@@
Provide your step-by-step reasoning first, and then print “The answer is (X)” where X is the answer choice (one capital letter), at the end of your response.
@@
Provide your step-by-step reasoning first, and then print “The answer is (X)” where X is the answer choice (one capital letter), at the end of your response.
Appendix M Symbolic-MoE as a Sparse Mixture-of-Expert